### v3

# Unlabeled Data or Pre-trained Model: Rethinking Semi-Supervised Learning and Pretrain-Finetuning

Abstract
Semi-supervised learning (SSL) alleviates the cost of data labeling process by exploiting unlabeled data, and has achieved promising results on various tasks such as image classification. Meanwhile, the Pretrain-Finetuning paradigm has garnered significant attention in recent years, and exploiting pre-trained models could also reduce the requirement of labeled data in downstream tasks. Therefore, a question naturally occurs: When the labeled data is scarce in the target tasks, should we exploit unlabeled data or pre-trained models? To answer this question, we select pre-trained Vision-Language Models (VLMs) as representative pretrain-finetuning instances and propose Few-shot SSL—a framework that enables fair comparison between these two paradigms by controlling the amount of labeled data used. Extensive experiments across various settings demonstrate that pre-trained VLMs generally outperform SSL methods in nearly all cases, except when the data has low resolution or lacks clear semantic structure. Therefore, we encourage future SSL research to compare with pre-trained models and explore deeper integration, such as using pre-trained knowledge to enhance pseudo-labeling. To support future research, we release our unified reproduction and evaluation framework. Codes are available here.
1 Introduction
Deep neural networks have demonstrated performance comparable to that of humans in certain supervised learning tasks, such as image classification [13, 20, 28]. However, these impressive results rely heavily on large amounts of labeled training data. In many real-world applications, obtaining such labeled data often demands substantial human and financial resources [54]. To reduce the dependence on labeled data and thereby cut costs, researchers have proposed Semi-Supervised Learning (SSL)—an effective approach that leverages unlabeled data to enhance model performance.
Recently, the Pretrain-Finetuning paradigm has gained significant attention, as models pre-trained on large-scale, readily available web data exhibit a strong ability to rapidly adapt to downstream tasks through fine-tuning. In image classification, for example, the semantic alignment acquired during pre-training enables models—such as CLIP [43], a pre-trained vision-language model—to perform zero-shot classification without task-specific supervision. Furthermore, building upon the rich knowledge gained during pre-training, various fine-tuning strategies (e.g., prompt tuning) have been proposed to boost task-specific performance with minimal labeled data [24, 25, 56, 57].
Both SSL and pretrain-finetuning paradigms are well-suited for downstream tasks with limited labeled data, leveraging either unlabeled data or pre-trained knowledge, respectively. This makes us come up with a question naturally:
When labeled data is scarce in the target task, should we rely on additional unlabeled data or pre-trained models? What are the respective advantages and disadvantages of each paradigm?
To answer this question, we propose Few-shot SSL, an evaluation framework designed to provide a fair comparison between SSL methods and pre-trained models under consistent conditions. This framework ensures a straightforward and intuitive comparison by aligning the number of labeled samples per class used for fine-tuning with the minimum number observed across various SSL settings. We adopt Vision-Language Models (VLMs), such as CLIP and its enhanced variants, as representative pre-trained models due to their strong generalization ability and proven effectiveness in zero- and few-shot learning across diverse domains. Both paradigms are then evaluated across four widely adopted semi-supervised settings within the unified Few-shot SSL framework, providing a comprehensive analysis of their respective advantages and limitations.
Experimental results indicate that SSL methods are well-suited for handling low-resolution datasets with limited semantic information. In contrast, pretrain-finetuning paradigm, without relying on unlabeled data, achieves competitive performance across various datasets and open-environment settings. The key takeaways under different settings are as follows:
  • Pre-trained models demonstrate strong generalization capabilities and high data efficiency, enabling them to outperform traditional SSL methods with only a small amount of labeled data. However, achieving optimal performance with pre-trained VLMs requires higher-resolution input images, typically exceeding 96$\times$96.
  • When the unlabeled data contains out-of-distribution classes, pre-trained VLMs can be leveraged to avoid unlabeled training, capitalizing on their extensive pre-trained knowledge and rapid adaptation capabilities. Additionally, pre-trained VLMs offer advantages such as robust result stability and high performance.
  • In the open-world setting, SSL methods perform better on low-resolution datasets, while pre-trained VLMs demonstrate balanced performance on both base and new classes in high-resolution datasets. Combined with our open-set analysis, we conclude that the pretrain-finetuning paradigm is better suited for scenarios where unlabeled data may contain previously unseen classes.
  • The low data requirements of pre-trained models enable them to effectively handle long-tailed distributions, offering a promising approach to addressing the inherent imbalance in many real-world datasets. Meanwhile, SSL methods remain well-suited for tasks characterized by less distinct semantic boundaries between categories, where learning directly from data representations proves advantageous.
In summary, the pretrain-finetune paradigm demonstrates strong performance across most tasks, except in scenarios with limited image resolutions or insufficient semantic information. To further advance research in related areas, we propose the following directions: (I) Incorporate comparisons with pre-trained models (e.g., VLMs) in SSL studies focused on image classification task. (II) Enhance the robustness of pre-trained models to low-resolution visual inputs, thereby improving their versatility across diverse real-world scenarios. (III) Explore methods for integrating semi-supervised learning with pre-trained model adaptation to further boost performance under limited supervision.
The remainder of this paper is structured as follows: In Section 2, we review the related work on the two paradigms, followed by a detailed explanation of their underlying mechanisms and our proposed Few-shot SSL framework in Section 3. In Section 4, we present extensive experiments to thoroughly investigate the performance of both paradigms under various settings. In Section 5, we provide an in-depth discussion on potential future directions, including the integration of these two paradigms and the impact of long-tailed data distributions on pretrain-finetuning approaches. Finally, in Section 6, we conclude our paper and offer recommendations for future research.
2 Related Work
Semi-Supervised Learning. The goal of Semi-Supervised Learning is to reduce the cost of data annotation by leveraging unlabeled data, under the assumption that both labeled and unlabeled samples belong to a predefined set of known classes. The primary objective is to assign accurate ground-truth labels to unlabeled instances. Existing SSL methods can be broadly categorized into three main approaches: Entropy Minimization [14, 15], Consistency Regularization[30, 38, 46], and Hybrid or Holistic Methods [4, 5]. Specifically, EntMin [14] promotes confident (i.e., low-entropy) predictions on unlabeled samples by introducing an additional loss term that minimizes prediction uncertainty, regardless of class. Pseudo-Labeling[32] assigns pseudo-labels to unlabeled data based on the model’s own predictions throughout training. FixMatch [48] combines consistency regularization and confidence-based filtering by enforcing consistency between weakly and strongly augmented samples, while only using high-confidence pseudo-labels for training, resulting in significant performance gains with limited labeled data.
Pretrain-Finetuning. The recent progress in pre-trained models has garnered significant attention, particularly with the emergence of Large Language Models like GPT [1] and Vision Language Models like CLIP [43]. These models, pre-trained on vast amounts of data, exhibit remarkable adaptability to downstream tasks, leading to the widespread adoption of the pretrain-finetuning paradigm. In image classification tasks, VLMs [2, 43, 55] are pre-trained on large-scale text-image pairs, leveraging shared embeddings of textual and visual representations to enable zero-shot classification. While these pre-trained VLMs have captured broad, generalized representations of images and text, efficiently adapting them to downstream tasks remains a significant challenge. Inspired by the success of prompt tuning in NLP [21, 33], many researchers have proposed to adapt VLMs by learning the prompts in end-to-end training. As the pioneering work, CoOp [57] for the first time introduces learnable prompts to transfer the task-specific knowledge to VLMs. Building on this idea, MaPLe [24] and PromptSRC [25] extend prompt tuning to both visual and textual encoders, enabling joint vision-language adaptation. In Few-shot SSL, we select CoOp and PromptSRC as the primary baselines.
3 Preliminary
Figure 1: Pipelines of semi-supervised learning and pretrain-finetuning.
In this section, we first present an overview of the working principles behind SSL and the pretrain–finetuning paradigms, as illustrated in Figure 1, along with their respective strategies for adapting to downstream tasks. We then describe the four semi-supervised settings used in our experiments, both in their original form and as adapted within our Few-shot SSL framework.
3.1 Semi-Supervised Learning
SSL leverages both a small set of labeled examples $\mathcal{D}_{l}=\{(x_{i},y_{i})\}_{i=1}^{N_{l}}$ and a larger pool of unlabeled examples $\mathcal{D}_{u}=\{x_{j}\}_{j=1}^{N_{u}}$ during training stage (see Figure 1 (a)). Let $f_{\theta}$ denote the model with parameters $\theta$. SSL jointly minimizes a supervised loss on $\mathcal{D}_{l}$ and an unsupervised loss on $\mathcal{D}_{u}$:
$\mathcal{L}(\theta)=\frac{1}{N_{l}}\sum_{i=1}^{N_{l}}\ell\bigl{(}f_{\theta}(x_{i}),y_{i}\bigr{)}\;+\;\lambda\frac{1}{N_{u}}\sum_{j=1}^{N_{u}}\mathcal{L}_{u}\bigl{(}f_{\theta},x_{j}\bigr{)},$ (1)
Here, $\ell(\cdot,\cdot)$ denotes the cross-entropy loss. The first term represents the supervised loss, while the second term captures the unsupervised loss, with $\lambda>0$ serving as a balancing factor between them. Although SSL reduces the dependence on labeled data, this advantage often comes with significant computational overhead during training, owing to reliance on large-scale use of unlabeled data.
3.2 Pretrain-Fintuning
Pre-trained models typically involve two stages before downstream evaluation (see Figure 1(b)). Taking CLIP [43] as an example, it employs an image encoder $f$ and a text encoder $g$ to extract feature representations $f(I)$ and $g(T)$ from an input image $I$ and its corresponding text caption $T$, respectively. During pre-training, the model minimizes a contrastive loss over $M$ image-text pairs:
$\mathcal{L}_{\text{pre-train}}=\frac{1}{M}\sum_{i=1}^{M}\Bigl{[}-\log\frac{\exp(s_{ii})}{\sum_{j=1}^{M}\exp(s_{ij})}\Bigr{]}\;+\;\frac{1}{M}\sum_{j=1}^{M}\Bigl{[}-\log\frac{\exp(s_{jj})}{\sum_{i=1}^{M}\exp(s_{ij})}\Bigr{]}\,.$ (2)
where $s_{ij}$ is the similarity score between $f(I_{i})$ and $g(T_{j})$, and The first and second terms represent the image-to-text and text-to-image contrastive losses ($\mathcal{L}_{\mathrm{img}\to\mathrm{text}}\ and\ \mathcal{L}_{\mathrm{text}\to\mathrm{img}}$). After pre-training, CLIP exhibits strong zero-shot performance by virtue of its large-scale domain knowledge and cross-modal alignment.
To enable rapid adaptation to downstream tasks, VLMs [56, 57] fine-tune a set of learnable prompts $P=[t_{1},t_{2},\dots,t_{b}]$, while keeping both the image and text encoders frozen. Given class name embeddings $[c_{1},c_{2},\dots,c_{N}]$, each class representation is formed by concatenating $P$ with $c_{i}$, and the prediction is computed as:
$p(y=i\mid I)=\frac{\exp\left(\cos\left(f(I),g([P,c_{i}])\right)/\tau\right)}{\sum_{j=1}^{N}\exp\left(\cos\left(f(I),g([P,c_{j}])\right)/\tau\right)},$ (3)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation and $\tau$ is a temperature parameter. During fine-tuning, only $P$ is updated using cross-entropy loss, enabling task-specific adaptation.
3.3 Few-Shot SSL
This section introduces four common semi-supervised settings and explains how they can be reformulated as a unified few-shot setting for pre-trained VLMs, referred to as Few-shot SSL. The dataset $\mathcal{D}$ used in this setting consists of a small labeled subset $\mathcal{D}_{l}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N_{l}}$ and a larger unlabeled subset $\mathcal{D}_{u}=\{\mathbf{x}_{j}\}_{j=1}^{N_{u}}$, where the number of labeled samples is significantly smaller than the unlabeled samples $(N_{l}\ll N_{u})$. The following provides a overview of each setting, and more details are shown in Appendix A.
  • Standard Semi-Supervised Setting [48]. In this setting, the labels for both subsets belong to the same known class set $\mathcal{Y}^{l}$. For VLMs, which do not rely on unlabeled data for training, the Few-shots’ SSL framework reduces this setting to a standard few-shot scenario, provided the number of labeled samples remains the same.
  • Open-Set Semi-Supervised Setting [41]. In this setting, the unlabeled dataset $\mathcal{D}_{u}$ may include instances from unseen classes, i.e., $\mathcal{Y}^{l}\subset\mathcal{Y}^{u}$ with $\mathcal{Y}^{\text{new}}=\mathcal{Y}^{u}\setminus\mathcal{Y}^{l}$. Since VLMs do not utilize unlabeled data, the open-set setting in Few-shots’ SSL reduces to the semi-supervised setting.
  • Open-World Semi-Supervised Setting [39]. This setting builds upon the open-set setting, aiming to classify both known and new classes. For VLMs, pre-training endows them with zero-shot classification capability, allowing for direct classification of both known and new classes. Therefore, the Open-world semi-supervised setting is analogous to the commonly used Base-to-Novel setting [56] for VLMs.
  • Long-Tailed Semi-Supervised Setting [29]. This setting features a highly imbalanced class distribution, where head classes have many samples and tail classes have few. The imbalance ratio (IR) is defined as the sample count ratio between the largest and smallest classes. Since VLMs use a fixed number of samples per class in few-shot settings, the head and tail classes become consistent due to the same number of samples. Thus, the Few-shot SSL framework reduces this setting to a standard few-shot scenario by matching the number of shots to that of the tail classes, ensuring fair comparison.
4 Experiments
4.1 Baselines Selection
We selected FixMatch [48] (standard semi-supervised learning), DWD [3] (open-set semi-supervised learning), OwMatch [39] (open-world semi-supervised learning), and CCL [58] (long-tailed semi-supervised learning) as representative methods across different settings. For the pretrain–finetuning paradigm, we consider CLIP [43], a pre-trained VLM, along with its prompt tuning variants—CoOp [57], the first to introduce prompt tuning for CLIP, and PromptSRC [25], a state-of-the-art method. Including these methods in our comparison allows for a direct observation of the lower and upper bounds for pre-trained VLMs in semi-supervised settings, leading to a more comprehensive analysis.
4.2 SSL vs. VLM Finetuning in Semi-Supervised Setting
Table 1: Accuracy on CIFAR-10, CIFAR-100, and STL-10. The best performance for each dataset is in bold, while the top scores in semi-supervised setting are underlined.
Method CIFAR-10 CIFAR-100 STL-10
40 labels 250 labels 4000 labels 400 labels 2500 labels 10000 labels 1000 labels
$\Pi$-Model - 45.74 $\pm$ 3.97 85.99 $\pm$ 0.38 - 42.75 $\pm$ 0.48 62.12 $\pm$ 0.11 73.77 $\pm$ 0.82
Pseudo-Labeling - 50.22 $\pm$ 0.43 83.91 $\pm$ 0.28 - 42.62 $\pm$ 0.46 63.79 $\pm$ 0.19 72.01 $\pm$ 0.83
Mean Teacher - 67.68 $\pm$ 2.30 90.81 $\pm$ 0.19 - 46.09 $\pm$ 0.57 64.17 $\pm$ 0.24 78.57 $\pm$ 2.39
MixMatch 52.46 $\pm$ 11.50 88.95 $\pm$ 0.86 93.58 $\pm$ 0.10 32.39 $\pm$ 1.32 60.06 $\pm$ 0.37 71.69 $\pm$ 0.33 89.59 $\pm$ 0.61
UDA 70.95 $\pm$ 5.93 91.18 $\pm$ 1.08 95.12 $\pm$ 0.18 40.72 $\pm$ 0.88 66.87 $\pm$ 0.22 75.50 $\pm$ 0.25 92.34 $\pm$ 0.56
ReMixMatch 80.90 $\pm$ 9.64 94.56 $\pm$ 0.05 95.28 $\pm$ 0.13 55.72 $\pm$ 2.06 72.57 $\pm$ 0.31 76.97 $\pm$ 0.56 94.77 $\pm$ 0.45
FixMatch (RA) 86.19 $\pm$ 3.37 94.93 $\pm$ 0.65 95.74 $\pm$ 0.05 51.15 $\pm$ 1.75 71.71 $\pm$ 0.11 77.40 $\pm$ 0.12 92.02 $\pm$ 1.50
FixMatch (CTA) 88.61 $\pm$ 3.35 94.93 $\pm$ 0.33 95.69 $\pm$ 0.15 50.05 $\pm$ 3.01 71.36 $\pm$ 0.24 76.82 $\pm$ 0.11 94.83 $\pm$ 0.63
ZSCLIP 79.30 $\pm$ 0.00 46.00 $\pm$ 0.00 97.40 $\pm$ 0.00
CoOp 80.60 $\pm$ 1.79 85.10 $\pm$ 0.29 89.13 $\pm$ 0.12 49.00 $\pm$ 0.83 57.43 $\pm$ 0.25 61.40 $\pm$ 0.14 98.23 $\pm$ 0.17
PromptSRC 86.03 $\pm$ 0.48 90.10 $\pm$ 0.08 93.90 $\pm$ 0.14 60.33 $\pm$ 0.17 66.33 $\pm$ 0.29 70.10 $\pm$ 0.22 98.57 $\pm$ 0.12
Table 2: Performance of SSL and pre-trained VLMs on the ImageNet and STL-10 datasets.
(a) Performance on the ImageNet dataset.
Method Accuracy
UDA [53] 68.78 $\pm$ 0.43 (10% labels)
FixMatch [48] 71.46 $\pm$ 0.52 (10% labels)
ZSCLIP [43] 66.70 $\pm$ 0.00 (0 labels)
CoOp [57] 73.43 $\pm$ 0.12 (2% labels)
PromptSRC [25] 74.03 $\pm$ 0.09 (2% labels)
(b) Performance on STL-10 at different image resolutions.
Resolution Method
ZSCLIP CoOp PromptSRC
16 $\times$ 16 49.30 $\pm$ 0.00 64.83 $\pm$ 0.45 76.97 $\pm$ 0.45
32 $\times$ 32 83.40 $\pm$ 0.00 87.63 $\pm$ 0.17 92.30 $\pm$ 0.36
64 $\times$ 64 95.00 $\pm$ 0.00 96.73 $\pm$ 0.12 97.67 $\pm$ 0.05
96 $\times$ 96 97.40 $\pm$ 0.00 98.23 $\pm$ 0.17 98.57 $\pm$ 0.12
Datasets and Experimental Setup. We evaluate both SSL and pre-trained VLMs on CIFAR-10/100 [27], STL-10 [9], and ImageNet [10], under varying annotation levels and augmentation strategies. For SSL methods (e.g., UDA [53], MixMatch [5]), we report the standard results as presented in FixMatch [48]. For VLMs, we simulate few-shot learning by training on the same labeled data only, without accessing unlabeled samples. All experiments on VLMs are conducted using the ViT-B/16 backbone and evaluated over three random seeds. Full implementation details are provided in Appendix A.2.
Experiment Results. As shown in Table 1, SSL methods generally perform well on lower-resolution datasets such as CIFAR-10/100 (32 $\times$ 32), surpassing zero-shot CLIP and its fine-tuning variants. This advantage stems from the difficulty pre-trained VLMs face in extracting meaningful features from lower-resolution inputs. In contrast, pre-trained VLMs excels on higher-resolution datasets like STL-10 (96 $\times$ 96), where even zero-shot CLIP surpasses the best-performing SSL baselines. Moreover, as shown in Table 2(a), on the ImageNet dataset, FixMatch requires a significantly larger proportion of labeled data (around 10%) but still lags behind in accuracy. By comparison, VLMs achieve over 73% accuracy using only 2% of labeled samples, with performance improving as more labeled data becomes available. In addition to accuracy, SSL methods typically involve much longer training times. For instance, training SSL models on CIFAR datasets takes several hours [4, 48], whereas prompt tuning approaches for VLMs [24, 25, 57] can be completed in just a few minutes. These observations highlight the complementary strengths of SSL and pretrain-finetuning paradigms, with the latter generally offering a higher performance ceiling and faster training speed.
Analysis. To investigate how input resolution affects pre-trained models’ performance, we evaluate several methods on the STL-10 dataset at varying resolutions (see Table 2(b)). As resolution increases, all VLMs exhibit substantial accuracy gains, highlighting their sensitivity to image resolution. We also experiment with different visual encoder backbones (presented in Appendix B) and find that ViT-based architectures consistently outperform ResNet-50 across all resolutions, indicating that the performance degradation is not due to an intrinsic limitation of Vision Transformer [11]. In fact, it stems from the fixed input size requirement (e.g., 224 $\times$ 224) of VLMs: low-resolution images must be upsampled via interpolation to satisfy this requirement, a process that inevitably introduces distortions and impairs the models’ ability to extract reliable visual features.
Takeaway 1: Pre-trained models demonstrate strong generalization capabilities and high data efficiency, enabling them to outperform traditional SSL methods with only a small amount of labeled data. However, achieving optimal performance with pre-trained VLMs requires higher-resolution input images, typically exceeding 96$\times$96.
4.3 SSL vs. VLM Finetuning in Open-Set Semi-Supervised Setting
Table 3: Performance comparison over three datasets. We report the mean accuracy averaged over three seeds, along with standard error.
Dataset FixMatch MPL IOMatch DWD-SL ZSCLIP CoOp PromptSRC
CIFAR-10/100 78.91 $\pm$ 0.15 70.95 $\pm$ 0.34 77.66 $\pm$ 0.22 80.05 $\pm$ 0.14 79.30 $\pm$ 0.00 87.87 $\pm$ 0.26 92.27 $\pm$ 0.12
ImageNet-30 70.07 $\pm$ 0.26 72.65 $\pm$ 0.70 79.23 $\pm$ 0.29 82.20 $\pm$ 0.38 98.80 $\pm$ 0.00 99.33 $\pm$ 0.12 99.30 $\pm$ 0.00
ImageNet-100 65.11 $\pm$ 0.32 68.43 $\pm$ 0.33 66.85 $\pm$ 0.19 82.81 $\pm$ 0.31 87.70 $\pm$ 0.00 91.37 $\pm$ 0.12 91.70 $\pm$ 0.08
Datasets and Experimental Setup. We evaluate both paradigms on standard open-set benchmarks, including CIFAR-10/100 [27], ImageNet-30 [22], and ImageNet-100 [6]. For CIFAR-10/100, we use 100 labeled images per class from CIFAR-10 as the supervised data and the entire CIFAR-100 dataset as the unsupervised data. For ImageNet-30 and ImageNet-100, we adopt a 50-shot setting, using 50 labeled samples per class from the selected in-distribution (ID) classes. It is important to note that the test set in this setting contains only those classes present in the training data. For the pretrain–finetuning paradigm, the experimental setup and VLM performance remain consistent with those described in the semi-supervised setting. Further details can be found in Appendix A.3.
Experiment Results. As shown in Table 3, zero-shot CLIP and its fine-tuned variants consistently outperform other methods across all datasets, with particularly strong results on ImageNet-30 and ImageNet-100. This advantage stems from their data efficiency—without reliance on noisy unlabeled data, VLMs are less impacted by out-of-distribution (OOD) samples in low-quality datasets. In contrast, SSL methods are more sensitive to such data quality issues, as they require specialized mechanisms to identify and filter out OOD samples. This reliance on additional OOD handling procedures inevitably affects both training efficiency and final performance, often leading to suboptimal results. Moreover, unlike SSL paradigms, pretrain–finetuning approaches benefit from extensive large-scale pretraining followed by targeted fine-tuning, resulting in consistently lower performance variance and a higher ceiling that scales with the amount of labeled data. Based on the above analysis, we conclude that the pretrain–finetuning paradigm is more suitable in scenarios where unlabeled data contains previously unseen classes.
Takeaway 2: When the unlabeled data contains out-of-distribution classes, pre-trained VLMs can be leveraged to avoid unlabeled training, capitalizing on their extensive pre-trained knowledge and rapid adaptation capabilities. Additionally, pre-trained VLMs offer advantages such as robust result stability and high performance.
4.4 SSL vs. VLM Finetuning in Open-World Semi-Supervised Setting
Table 4: Average accuracy on CIFAR-10/100 and ImageNet-100 datasets with both new class ratio and label ratio of 50%.
Method CIFAR-10 CIFAR-100 ImageNet-100
Base New All Base New All Base New All
FixMatch [48] 71.5 50.4 49.5 39.6 23.5 20.3 65.8 36.7 34.9
DS3[17] 77.6 45.3 40.2 55.1 23.7 24.0 71.2 32.5 30.8
CGDL [49] 72.3 44.6 39.7 49.3 22.5 23.6 67.3 33.8 31.9
DTC [19] 53.9 39.5 38.3 31.3 22.9 18.3 25.6 20.8 21.3
RankStats [18] 86.6 81.0 82.9 36.4 28.4 23.1 47.3 28.7 40.3
SimCLR [8] 58.3 63.4 51.7 28.6 21.1 22.3 39.5 35.7 36.9
ORCA [7] 88.2 90.4 89.7 66.9 43.0 48.1 89.1 72.1 77.8
NACH [16] 89.5 92.2 91.3 68.7 47.0 52.1 91.0 75.5 79.6
OpenLDN [44] 95.7 95.1 95.4 73.5 46.8 60.1 89.6 68.6 79.1
OpenCon [50] 89.3 91.1 90.4 69.1 47.8 52.7 90.6 80.8 83.8
OwMatch [39] 93.0 95.9 94.4 74.5 55.9 65.1 91.7 72.0 81.8
OwMatch+ [39] 96.5 97.1 96.8 80.1 63.9 71.9 91.5 79.6 85.5
ZSCLIP [43] 88.0 88.1 79.3 57.4 53.1 46.0 89.5 87.9 87.7
CoOp [57] 91.1 90.4 81.3 65.4 51.0 48.4 89.7 83.2 85.5
PromptSRC [25] 93.4 91.1 83.9 73.0 63.2 59.7 91.7 89.5 89.5
Datasets and Experimental Setup. We conduct evaluations on CIFAR-10/100 [27], and ImageNet-100 [6]. In each dataset, the classes are divided into known (base) and new classes, with the proportion of new classes set to 50% by default. A portion of the samples from the known classes are labeled according to a given label ratio, while the remaining part, together with all the new class data, formed the unlabeled set. This setup requires the model to simultaneously distinguish and classify both base and new classes. Within the context of pretrain-finetuning paradigm, this is commonly known as the Base-to-New setting. For more detailed settings, please refer to Appendix A.4.
Experiment Results. In the open-world setting, SSL methods aim to identify unseen classes from unlabeled data during training, typically using clustering-based pseudo-labeling, while evaluating both base and novel classes at test time. As shown in Table 4, these methods achieve the best performance on CIFAR-10 and CIFAR-100, further supporting the conclusions drawn from the semi-supervised setting analysis. In contrast, on the ImageNet-100 dataset, zero-shot CLIP already approaches the performance of OwMatch on base classes and significantly outperforms it on novel classes. This can be attributed to the richer feature representations captured by pre-trained models as image resolution increases, enabling them to better leverage prior knowledge. Notably, pretrain-finetuning approaches exhibit a smaller performance gap between base and novel classes on ImageNet-100 compared to SSL methods, highlighting their strong generalization capabilities, which allow for rapid adaptation to new classes.
Takeaway 3: In the open-world setting, SSL methods perform better on low-resolution datasets, while pre-trained VLMs demonstrate balanced performance on both base and new classes in high-resolution datasets. Combined with our open-set analysis, we conclude that the pretrain-finetuning paradigm is better suited for scenarios where unlabeled data may contain previously unseen classes.
4.5 SSL vs. VLM Finetuning in Long-Tailed Semi-Supervised Setting
Table 5: Test accuracy of SSL and pretrained VLMs under consistent imbalance ratio setting on CIFAR10-LT and CIFAR100-LT datasets.
Method CIFAR10-LT CIFAR100-LT
$\gamma=\gamma_{l}=\gamma_{u}=100$ $\gamma=\gamma_{l}=\gamma_{u}150$ $\gamma=\gamma_{l}=\gamma_{u}=10$ $\gamma=\gamma_{l}=\gamma_{u}=20$
$N_{1}=500$ $M_{1}=4000$ $N_{1}=1500$ $M_{1}=3000$ $N_{1}=500$ $M_{1}=4000$ $N_{1}=1500$ $M_{1}=3000$ $N_{1}=50$ $M_{1}=400$ $N_{1}=150$ $M_{1}=300$ $N_{1}=50$ $M_{1}=400$ $N_{1}=150$ $M_{1}=300$
Supervised 47.3 $\pm$ 0.95 61.9 $\pm$ 0.41 44.2 $\pm$ 0.33 58.2 $\pm$ 0.29 29.6 $\pm$ 0.57 46.9 $\pm$ 0.22 25.1 $\pm$ 1.14 41.2 $\pm$ 0.15
w/ LA [37] 53.3 $\pm$ 0.44 70.6 $\pm$ 0.21 49.5 $\pm$ 0.40 67.1 $\pm$ 0.78 30.2 $\pm$ 0.44 48.7 $\pm$ 0.89 26.5 $\pm$ 1.31 44.1 $\pm$ 0.42
FixMatch + LA [37] 75.3 $\pm$ 2.45 82.0 $\pm$ 0.36 67.0 $\pm$ 2.49 78.0 $\pm$ 0.91 47.3 $\pm$ 0.58 58.6 $\pm$ 0.36 41.4 $\pm$ 0.93 53.4 $\pm$ 0.32
w/ DARP [26] 76.6 $\pm$ 0.92 80.8 $\pm$ 0.62 68.2 $\pm$ 0.24 76.7 $\pm$ 1.35 50.4 $\pm$ 0.65 59.0 $\pm$ 0.53 43.6 $\pm$ 1.05 55.3 $\pm$ 0.84
w/ CReST+ [51] 76.7 $\pm$ 1.31 81.1 $\pm$ 0.57 70.9 $\pm$ 1.14 77.9 $\pm$ 0.71 44.0 $\pm$ 0.66 57.1 $\pm$ 1.10 40.6 $\pm$ 0.52 52.3 $\pm$ 0.30
w/ DASO [40] 77.9 $\pm$ 0.88 82.5 $\pm$ 0.60 71.0 $\pm$ 1.68 79.0 $\pm$ 2.23 50.1 $\pm$ 0.75 61.0 $\pm$ 0.72 43.5 $\pm$ 0.81 56.1 $\pm$ 0.72
FixMatch + ACR [52] 81.6 $\pm$ 0.19 84.1 $\pm$ 0.20 77.0 $\pm$ 1.19 82.0 $\pm$ 0.41 51.1 $\pm$ 0.80 61.3 $\pm$ 0.26 46.1 $\pm$ 0.34 56.5 $\pm$ 0.33
FixMatch + CPE [35] 80.7 $\pm$ 0.96 84.4 $\pm$ 0.29 76.8 $\pm$ 0.53 82.3 $\pm$ 0.36 50.6 $\pm$ 0.45 59.8 $\pm$ 0.13 43.6 $\pm$ 0.88 55.8 $\pm$ 0.65
FixMatch + CCL [58] 84.5 $\pm$ 0.38 86.2 $\pm$ 0.35 81.5 $\pm$ 0.99 84.0 $\pm$ 0.21 53.5 $\pm$ 0.49 63.5 $\pm$ 0.39 46.8 $\pm$ 0.45 57.5 $\pm$ 0.55
ZSCLIP [43] 79.30 $\pm$ 0.00 46.00 $\pm$ 0.00
CoOp [57] 80.73 $\pm$ 1.90 84.20 $\pm$ 0.50 80.60 $\pm$ 1.79 82.23 $\pm$ 0.75 49.83 $\pm$ 1.04 55.23 $\pm$ 0.50 44.83 $\pm$ 0.45 52.57 $\pm$ 0.12
PromptSRC [25] 86.50 $\pm$ 0.37 88.70 $\pm$ 0.08 86.03 $\pm$ 0.48 87.67 $\pm$ 0.42 60.33 $\pm$ 0.17 64.37 $\pm$ 0.34 58.03 $\pm$ 0.46 62.83 $\pm$ 0.48
Table 6: Test accuracy under inconsistent imbalance ratio setting ($\gamma_{l}\neq\gamma_{u}$) on STL10-LT and ImageNet-127 datasets. $\gamma=10/20$ for STL10-LT dataset.
Method STL10-LT STL10-LT ImageNet-127
$\gamma=10$ $\gamma=20$ $\gamma=\text{None}$
$N_{1}=150$ $M=100k$ $N_{1}=450$ $M=100k$ $N_{1}=150$ $M=100k$ $N_{1}=450$ $M=100k$ 32 $\times$ 32 64 $\times$ 64
FixMatch 56.1 $\pm$ 2.32 72.4 $\pm$ 0.71 47.6 $\pm$ 4.87 64.0 $\pm$ 0.27 29.7 42.3
   w/ DARP [26] 66.9 $\pm$ 1.66 75.4 $\pm$ 1.05 59.9 $\pm$ 2.17 72.3 $\pm$ 0.60 30.5 42.5
   w/ CReST+ [51] 56.0 $\pm$ 3.13 71.5 $\pm$ 0.96 55.4 $\pm$ 3.03 68.5 $\pm$ 1.15 32.5 44.7
   w/ DASO [40] 61.3 $\pm$ 1.84 78.4 $\pm$ 0.96 68.5 $\pm$ 1.78 75.3 $\pm$ 0.44 40.9 55.9
   w/ CCL [58] 79.1 $\pm$ 0.43 84.8 $\pm$ 1.05 77.1 $\pm$ 0.33 83.1 $\pm$ 0.18 61.5 67.8
ZSCLIP [43] 97.4 $\pm$ 0.00 46.00 $\pm$ 0.00
CoOp [57] 97.77 $\pm$ 0.19 98.13 $\pm$ 0.17 97.73 $\pm$ 0.25 97.83 $\pm$ 0.09 49.83 $\pm$ 1.04 55.23 $\pm$ 0.50
PromptSRC [25] 98.27 $\pm$ 0.05 98.47 $\pm$ 0.05 98.13 $\pm$ 0.12 98.30 $\pm$ 0.08 60.33 $\pm$ 0.17 64.37 $\pm$ 0.34
Datasets and Experimental Settings. We compare different methods on four long-tailed datasets: CIFAR-10-LT [27], CIFAR-100-LT [27], STL-10-LT [9], and ImageNet-127 [12]. For synthetic long-tailed datasets (CIFAR, STL-10), the class-wise labeled/unlabeled sample counts follow an exponential decay controlled by imbalance ratios $\gamma_{l}$ and $\gamma_{u}$ [52]. For ImageNet-127, a naturally long-tailed dataset, we follow [58] and use 10% of the training data as labeled. To ensure fair comparison, we adopt the same splits and protocols as SSL methods [39], and report performance of VLMs in a few-shot setting where the number of shots is determined by the smallest labeled class (e.g., 5-shot for CIFAR-10 when $\gamma_{l}=100$ and $N_{1}=500$). Full setup details and shot derivations are available in Appendix A.5.
Experiment Results. As shown in Table 5, in the long-tailed semi-supervised setting, SSL methods are designed to ensure that the learning process effectively captures all classes, even when the distribution is dominated by a few head classes. However, as shown in Table 1, their performance often declines in the same datasets compared to standard SSL, highlighting their ongoing vulnerability to the challenges of long-tailed distributions. In contrast, the evaluation results of pre-trained VLMs underscore their efficiency, which can be fine-tuned with only a small amount of labeled data and still perform well, even on low-resolution datasets, a common challenge for these models. As shown in Table 6, pre-trained models perform well on highly imbalanced datasets like STL-10 but struggle on semantically coarse datasets such as ImageNet-127 [12]. ImageNet-127 is obtained by collapsing the 1,000 fine-grained ImageNet-1K classes into 127 coarse-grained categories according to their positions in the WordNet semantic hierarchy, yielding a long-tailed distribution (e.g., grouping “abacus,” “acoustic guitar,” and “analog clock” under the broader category “device”). Zero-shot CLIP evaluations reveal that this hierarchical coarsening injects semantic ambiguity, making it challenging for pretrained VLMs to confidently associate a fine-grained concept with a specific image. Although downstream fine-tuning can partially mitigate this issue, the underlying semantic constraints remain. Detailed analysis and visualization are shown in Appendix C.
Takeaway 4: The low data requirements of pre-trained models enable them to effectively handle long-tailed distributions, offering a promising approach to addressing the inherent imbalance in many real-world datasets. Meanwhile, SSL methods remain well-suited for tasks characterized by less distinct semantic boundaries between categories, where learning directly from data representations proves advantageous.
5 Discussion
Potential for combining SSL and pre-trained models. Integrating SSL with pre-trained models offers new opportunities for semi-supervised classification. While SSL leverages unlabeled data to reduce labeling costs, pre-trained models excel in zero-shot learning and generalization. Their combination can effectively address challenges such as few-shot learning, class imbalance, and other specialized scenarios. For instance, PromptKD [34] introduces a two-stage, prompt-driven knowledge distillation framework that aligns teacher and student models through learnable prompts, achieving effective knowledge transfer using a diverse mix of ID and OOD unlabeled images, resulting in significant improvements across various datasets. Similarly, UPL [23] automatically generates high-quality prompts from unlabeled data, reducing the dependence on large-scale labeled datasets and improving overall model fit. GRIP [36] represents another promising direction, proposing a pseudo-label-based fine-tuning method that effectively combines SSL principles with VLMs. Clearly, leveraging pseudo-labeled data, semi-supervised learning concepts, and the generalization strengths of VLMs offers a promising path for enhancing performance across a wide range of downstream tasks, prompting a fresh perspective on hybrid learning strategies.
Effects of training with long-tailed data on pre-trained models. When training data follows a long-tailed distribution, excessive fine-tuning of pre-trained models can be detrimental, particularly to the performance on tail classes. As highlighted in [47], aggressive fine-tuning can disrupt the class-conditional structure learned during pre-training, resulting in poor representation for underrepresented classes. To address this issue, they proposed LIFT, a lightweight fine-tuning approach that modifies only a small portion of the model (e.g., prompts or adapter layers), effectively adapting to long-tailed distributions while preserving the semantic consistency of the pre-trained model. This demonstrates that leveraging pre-trained models can help mitigate the adverse effects of long-tailed distributions. However, current approaches primarily focus on reducing fine-tuning intensity, often at the cost of head class performance. This highlights the need for strategies that can guide models to fully adapt to imbalanced data while minimizing learning bias.
6 Conclusion and Recommendation
Our experiments provide a comprehensive and fair analysis of the strengths and weaknesses of SSL and pretrain-finetuning paradigms. When should each approach be preferred in real-world applications, and what are their advantages and limitations? Our key insights are as follows:
  • The SSL Paradigm: Best suited for scenarios with low-resolution images or ambiguous class semantics. However, it struggles with noisy, unlabeled data, often requiring more training time and a larger number of training parameters.
  • The Pretrain-Finetuning Paradigm: Ideal for handling imbalanced data distributions and noisy, unlabeled datasets, but tends to underperform on low-resolution images due to its heavy reliance on high-resolution visual inputs and clear semantic information.
Our study also points to promising directions for future research, including but not limited to designing hybrid approaches that effectively combine the strengths of both paradigms:
  • Datasets: Higher-resolution and more challenging datasets should be included to better evaluate the performance of semi-supervised algorithms and pre-trained models, in line with the continued advancement of image resolution.
  • The SSL Paradigm: Future work on semi-supervised learning for image classification should include comparisons with pre-trained models (e.g., VLMs) under matched labeled data conditions.
  • The Pretrain-Finetuning Paradigm: Pre-trained models should be made more robust to low-resolution images to enhance their generalization across a wider range of real-world scenarios.
  • General Directions: Explore methods that integrate semi-supervised learning with pre-trained model adaptation to further improve their performance in labeled data-scarce environments.
One limitation of our study is that we focused solely on evaluating CLIP-based pre-trained models for image classification. In the future, we will conduct more extensive comparisons, hoping that our results and findings will inspire further advancements in SSL and Pretrain-Finetuning paradigms.
Appendix A Detailed Experimental Settings
A.1 Overal Experimental Settings
The overall VLM setup experiments are all conducted on a single NVIDIA A800 GPU, and the SSL method directly reported the results of the most recognized articles.
A.2 Detailed Semi-supervised Datasets and Experimental Setup
Datasets Setup. We evaluate the performance of both classical SSL methods and VLMs on standard image classification benchmarks: CIFAR-10, CIFAR-100 [27], STL-10 [9], and ImageNet [10]. Each dataset is tested under multiple supervision levels, typically using 40, 250, or 4000 labeled samples depending on the benchmark. We also consider different augmentation strategies (e.g., weak/strong) following the standard SSL protocol.
Experimental Setup. For classical SSL baselines (FixMatch [48], MixMatch [5], ReMixMatch [4]), models are pre-trained on the labeled subset and optimized using a combination of supervised and consistency-based unsupervised losses. We directly use results reported in the FixMatch paper to maintain consistency. The unlabeled data is drawn from the remaining training set without labels. For Vision-Language Models (VLMs), we adopt the ViT-B/16 architecture across all experiments. Since VLMs are pre-trained on large-scale image-text pairs, we do not require any additional unlabeled samples. Instead, we fine-tune them using only the labeled portion of the data, treating the task as a pure few-shot classification problem. We use prompt tuning methods such as CoOp and PromptSRC (depending on the experiment), and all models are trained for 50 epochs using the labeled data only. Hyperparameters are kept consistent with those used in the Base-to-New setting. For CoOp, we follow the settings reported in the original paper: 4-shot training for 100 epochs and 25/100/400-shot training for 200 epochs. PromptSRC strictly follows the configurations reported in its paper, with key parameters as follows: both textual and visual prompts consist of 4 tokens, and the first nine transformer layers are modified with learnable prompt tokens. All PromptSRC experiments are run for 20 epochs. Experiments are conducted on a single NVIDIA A800 GPU.
A.3 Detailed Open-set Semi-supervised Datasets and Experimental Setup
Dataset Setup. We follow commonly used OSSL benchmarks, including CIFAR-10/100, ImageNet-30, and ImageNet-100. A detailed description of these datasets is provided below:
  • CIFAR-10/100. This task uses CIFAR-10 as the labeled dataset and CIFAR-100 as the unlabeled dataset. While the entire CIFAR-100 dataset serves as the unlabeled portion, we sample 100 images per class from CIFAR-10 to simulate a scarce-label scenario. For our VLMs, since training is still conducted solely on the labeled CIFAR-10 data in a few-shot manner, the results reflect a 100-shot training setup for CIFAR-10 under the SSL setting.
  • ImageNet-30. This task uses the ImageNet-30 dataset [22], a 30-class subset of ImageNet. Following [42], we select 5% of the data (approximately 50 images per class) from the first 20 alphabetically sorted classes as labeled data, and treat the rest as unlabeled. For VLMs, the minimum number of samples among these 50 classes is 754; 10% of that equals roughly 70. To maintain consistency and simplify comparisons, we fix the labeled set to 50-shot per class for fine-tuning.
  • ImageNet-100. This task is based on the ImageNet-100 dataset, a 100-class subset of ImageNet as defined in [6]. The classes are split evenly into 50% in-distribution (ID) and 50% out-of-distribution (OOD) based on alphabetical order. For the labeled data, we select 50 images per class from the first 20 ID classes, while the remaining data from all 100 classes is used as unlabeled data. VLMs are fine-tuned using only the 50-shot labeled data from the first 20 ID classes, and evaluated on the corresponding test set following the same setup.
Experimental Setup. Since VLMs do not utilize unlabeled data, the Open-set semi-supervised setting on above datasets essentially becomes a few-shot learning problem with a reduced number of classes. The results of OSSL methods refer to DWD-SL [3].
A.4 Detailed Open-World Semi-Supervised Datasets and Experimental Setup
Dataset Setup. We evaluate our approach on CIFAR-10/100 [27], ImageNet-100 [6], and Tiny ImageNet [31]. Specifically, the ImageNet-100 dataset contains 100 classes sub-sampled from ImageNet-1k following the selection in [10]. For all datasets, we split classes into seen and new categories with a predefined new class ratio (defaulting to 50% unless otherwise specified). From the seen classes, a fixed ratio of data is randomly selected as labeled data, while the remaining seen samples, together with all samples from the new classes, are assigned to the unlabeled set.
Experimental Setup. Under the Base-to-New setting commonly used for evaluating VLMs [25, 57], each dataset’s classes are split evenly into base and new groups. VLMs are trained exclusively on the base classes and then evaluated separately on both base and new classes to assess their generalizability. OWSSL setting closely aligns with the Base-to-New setting, but differs in how unseen (new) classes are treated during training: SSL methods assume the unlabeled pool contains OOD (new) samples and employ clustering or OOD-detection mechanisms to identify and leverage them. While VLMs use only labeled base-class data, without any exposure to OOD examples. At evaluation time, VLMs simply rely on their pre-trained image–text alignment to generalize to new categories, whereas OWSSL methods benefit from the self-label strategy during training [39, 45].
We report OWSSL results using OwMatch [39]. For VLMs, we adopt ViT-B/16 backbones across all variants. CoOp and PromptSRC are fine-tuned using only the labeled data under varying label ratios (e.g., 1%, 5%, 10%), and are trained for 50 epochs following [25, 57]. All other hyperparameters remain consistent across datasets and methods unless otherwise stated.
A.5 Detailed Long-Tailed Datasets and Experimental Settings
Dataset Setup. We evaluate methods on widely-used long-tailed SSL benchmarks: CIFAR-10-LT, CIFAR-100-LT [27], STL-10-LT [9], and ImageNet-127 [12]. For synthetic datasets, the number of samples for each class $c$ is defined as $N_{c}=N_{1}\cdot\gamma_{l}^{-\frac{c-1}{C-1}}$, where $N_{1}$ is the number of samples in the most frequent class and $\gamma_{l}$ is the imbalance ratio. Unlabeled data follows a similar distribution with $\gamma_{u}$.
  • CIFAR-10-LT: Following DASO [52], we adopt $(N_{1},M_{1})\in\{(500,4000),(1500,3000)\}$ and $\gamma_{l}=\gamma_{u}\in\{100,150\}$. We also consider uniform/reversed unlabeled distributions with $\gamma_{u}\in\{1,1/100\}$ and $\gamma_{l}=100$ fixed.
    Based on the minimum labeled class size, we define the corresponding few-shot settings for VLMs:
    • 5 shots for $(\gamma=100,\ N_{1}=500)$
    • 15 shots for $(\gamma=100,\ N_{1}=1500)$
    • 4 shots for $(\gamma=150,\ N_{1}=500)$
    • 20 shots for $(\gamma=150,\ N_{1}=3000)$
  • CIFAR-100-LT: We adopt $(N_{1},M_{1})\in\{(50,400),(150,300)\}$ with $\gamma_{l}=\gamma_{u}\in\{10,20\}$. The corresponding shots for VLMs are:
    • 5 shots for $(\gamma=10,\ N_{1}=50)$
    • 15 shots for $(\gamma=10,\ N_{1}=150)$
    • 3 shots for $(\gamma=20,\ N_{1}=50)$
    • 8 shots for $(\gamma=20,\ N_{1}=150)$
  • STL-10-LT: We only control labeled imbalance ratio with $\gamma_{l}\in\{10,20\}$, as unlabeled data lacks class labels.
  • ImageNet-127: This is a real-world long-tailed dataset with naturally imbalanced label distribution. Following CCL [58], we sample 10% of training data for the labeled set. Due to this sampling strategy, there is no fixed $\gamma$ value.
Experimental Setup. Since the LTSSL setting only affects the training set distribution, we evaluate all models on the same balanced test sets as prior work. For example, in CIFAR-10-LT with $\gamma_{l}=\gamma_{u}=100$, $N_{1}=500$ implies that the rarest class has only 5 labeled samples. For VLMs, we treat this as a 5-shot setup and do not use any unlabeled data. All VLMs follow the same architecture (ViT-B/16), hyperparameters, and optimization schedules as in Section 4.2.
Appendix B Detailed CIFAR10/100 Results under Different Backbones
To ensure a consistent amount of training data and other training parameters, we trained PromptSRC with various visual encoders on the CIFAR dataset, running each configuration three times. As shown in Table 7, the experimental results demonstrate a gradual increase in model performance as the visual encoder’s parameter count grows. This suggests that the suboptimal performance on low-resolution images is not attributable to the visual encoder itself.
Table 7: Performance of different visual backbones on the CIFAR-10/100 datasets.
Dataset Backbones
ResNet-50 ViT-B/32 ViT-B/16
CIFAR10 70.23$\pm$0.56 82.00$\pm$0.70 85.10$\pm$0.29
CIFAR100 39.87$\pm$0.48 54.00$\pm$0.29 57.43$\pm$0.25
Appendix C Detailed Analysis and Visualization of ImageNet-127
ImageNet-127 compresses the original 1,000 fine-grained ImageNet classes into 127 higher-level categories, inevitably producing a long-tailed distribution. For example, 124 distinct classes—including abacus, acoustic guitar, and analog clock—are merged into the super-category device, whereas a single class such as bubble constitutes the entire globule category. Thus, categories that absorb many fine-grained classes become head classes, while those represented by only one class form the tail.
This hierarchical regrouping introduces substantial semantic ambiguity: the pre-training stage of VLMs fails to adequately encode the latent hierarchy, resulting in frequent misalignments, as shown in Figure 2(a). Although fine-tuning enables pre-trained VLMs to better bridge these granularity gaps and exploit limited downstream labels, the fundamental semantic imbalance remains, as further illustrated in Figure 2(b).
(a) Zero-shot alignment of fine-grained and coarse-grained category embeddings.
(b) Fine-tuned alignment of fine-grained and coarse-grained category embeddings.
Figure 2: Comparison of category embedding alignment in zero-shot and fine-tuned models. The left plot shows the misalignment between fine-grained and coarse-grained category embeddings before fine-tuning, while the right plot illustrates the improved alignment achieved through fine-tuning.